OpenStudy (anonymous):

If 9A =5(C-32) and A=100, find C.

OpenStudy (anonymous):

c-32=900/5, c=212.

OpenStudy (saifoo.khan):

9A =5(C-32) 9(100) = 5(C-32) 900/5 = C -32

OpenStudy (saifoo.khan):

C = 180 + 32 C = 212

OpenStudy (anonymous):

Aloha :)

OpenStudy (saifoo.khan):

[: